a) Define crystal and unit cell

b) Explain de Broglie wave and the photoelectric effect.

c) A wire of length 7 π‘š and radius 2.37 π‘šπ‘š is stretched by a load of 12370 π‘”π‘š and increased by 1.78% of the original length. Find the stress, strain, and Young’s modulus of the material of the wire.Β 


1
Expert's answer
2021-12-07T20:12:34-0500

a) Crystal is a complex system that obtains all mechanical properties of a solid. A unit cell is the smallest element of the crystal that has such properties.

b) De Broglie wavelength is a wavelength a body would be if it were a particle. The photoelectric effect: metals emit electrons when light of the corresponding wavelength is shone on them.

c) The strain: 1.78/100=0.0178.

The stress:


"\sigma=F/A=mg/(\pi r^2)=6.87\text{ MPa}."


The young's modulus:


"E=\frac\sigma\epsilon=386\text{ MPa}."